# Dynamic modes of active Potts models with factorizable numbers of states

### v1

@@ Abstract
We studied the long-term nonequilibrium dynamics of $q$-state Potts models with $q=4$, $5$, $6$, and $8$ using Monte Carlo simulations on a two-dimensional square lattice. When the contact energies between the nearest neighbors for the standard Potts models are used, cyclic changes in the $q$ homogeneous phases and $q$-state coexisting wave mode appear at low and high flipping energies, respectively, for all values of $q$. However, for a factorizable $q$ value, dynamic modes with skipping states emerge, depending on the contact energies. For $q=6$, a spiral wave mode with three domain types (one state dominant or two states mixed) and cyclic changes in three homogeneous phases are found. Although three states can coexist spatially under thermal equilibrium, the scaling exponents of the transitions to the wave modes are modified from the equilibrium values.
@@ I Introduction
Various types of pattern formation have been observed in nonequilibrium systems [1, 2, 3, 4, 5, 6, 7]. Classically, they have been called self-organizing systems [1, 2], dissipative structures [1], and synagetics [2, 3], and recently, nonreciprocal systems [8, 9, 10, 11], featuring different aspects. In biological systems, spatiotemporal patterns have been observed from microscopic to macroscopic scales (cells and tissues to animal populations) [3, 7, 12, 13, 14, 15]. Under far-from-equilibrium conditions, the thermal fluctuations occurring in macroscopic patterns are typically negligible and the dynamics can be explained using deterministic continuum equations. In contrast, the thermal fluctuations can be significant in microscopic patterns, such as intracellular waves. However, the effects of the fluctuations and nucleation are not understood so far.
@@
Lattice models, including Ising and Potts models, have been widely used to study the phase transitions occurring at thermal equilibrium [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. The nonequilibrium dynamics have been investigated by imposing periodic external fields [31, 32] and spatial temperature inhomogeneity [33]. Recently, the cyclic flipping energies of three or four states have been introduced to these lattice models, and their pattern formations have been studied [34, 35, 36, 37, 38, 39]. We have used the standard Potts models [16, 17] for three [34, 35] and four [36] states in a two-dimensional (2D) square lattice under both cyclic symmetry [34, 36] and asymmetry [35, 36]; Hence, the transitions between dynamic modes, such as the cyclic changes of homogeneous phases and spiral waves, were revealed. These dynamics were also obtained in an off-lattice model for undulating fluid membranes [40]. These models exhibit stable spatiotemporal patterns in the long-term limit, unlike lattice Lotka–Volterra models, which exhibit an absorbing transition to a uniform state [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Manacorda and Fodor reported wave and global oscillation modes using a lattice model, which allows the occupancy of multiple particles at each site [37]. Further, Avni et al. have added a nonreciprocal interaction to the Ashkin-Teller model with a pairwise spin interaction (it is also called four-state vector-Potts and clock models [16]) and revealed changes in the scaling exponents of the second-order transition between disordered and ordered phases [38, 39]. However, this ordered phase maintains only for a short period (tproportional-$t\propto$ system size $N$) starting from a uniform initial state.
@@
With the aim of clarifying the effects of thermal fluctuations on pattern formation and nonequilibrium phase transitions, this study investigates several transitions between the dynamic modes in the long-term limit ($t\to\infty$), which are independent of the initial states. We use general $q$-state Potts models with $q=4$, $5$, $6$, and $8$ under cyclically symmetric conditions. Various dynamic patterns are formed, particularly for a factorizable $q$ number. Although phases comprising factorized numbers of states have been reported at thermal equilibrium [21, 22, 23, 24, 55], to the best of our knowledge, spatiotemporal patterns with factorized symmetry have not been reported. We reveal several types of patterns with factorized symmetries and transitions between them.
@@
The model and methods are described in Sec. II. Simulation results are presented and discussed in Sec. III. The results for $q=6$, $5$, $4$, and $8$ are described in Secs. III.1,  III.2III.3, and III.4, respectively. The relationship with other models is discussed in Sec. IV. Finally, a summary is presented in Sec. V.
@@ II Active Potts Models
In standard Potts models, $J_{s_{i},s_{j}}=J_{0}\delta_{s_{i},s_{j}}$ is used. In equilibrium systems, each state can additionally have a self-energy $\varepsilon_{s}$, and the forward and backward flip rates is $\exp(-\Delta H_{s_{i}s^{\prime}_{i}})$ for flipping of a single site from $s$ to $s^{\prime}$. The thermal energy $k_{\mathrm{B}}T$ is normalized to unity, $\Delta H_{s_{i}s^{\prime}_{i}}=\Delta H_{\mathrm{int}}-h_{s,s^{\prime}}$, and the flipping energy $h_{s,s^{\prime}}=\varepsilon_{s}-\varepsilon_{s}^{\prime}$. The cyclic sum of the flipping energy is $\sum_{k=0}^{q}h_{k,[k+1]}=0$ at equilibrium, where $[k]$ indicates $k\bmod q$. We extended this model to a nonequilibrium situation, in which $\sum_{k=0}^{q}h_{k,[k+1]}>0$, but $h_{ss^{\prime}}=-h_{s^{\prime}s}$ remains [34]. Hence, the detailed balance can be locally satisfied for flips between $s$ and $s^{\prime}$, but not globally for cycles ($s=0\to 1\ ...\to q-1\to 0$). For $q=3$, this corresponds to the rock–paper–scissors relationship. Note that this situation can be realized by reactions on a catalytic surface [56, 57, 58, 59, 60] and molecular transport through a membrane [61, 62, 63, 64, 40]. In the context of surface reactions, the state $s=0$ represents an unoccupied site, $s=1$ is a reactant-bound state, and $s=2$ to $q-2$ and $s=q-1$ are intermediate and final products, respectively. The bulk reaction energy for the final product is expressed as $\sum_{k=0}^{q}h_{k,[k+1]}$. In the case of molecular transport, the molecules have $q-1$ intermediate states in the membrane.
@@
Figure 1: Six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$ and $L=128$. (a–d) Snapshots. The yellow, green, light-blue, dark-blue, magenta, and red sites (light to dark in grayscale) represent $s=0$, $1$, $2$, $3$, $4$, and $5$, respectively. (a) Homogeneous cycling of six states (HC6) at $h=0.8$ and $J_{k,[k+2]}=-1$. (b) Waves of six states (W6) at $h=1.5$ and $J_{k,[k+2]}=-1$. (c) Waves of three states (W3) at $h=1$ and $J_{k,[k+2]}=0.5$. (d) Mixing of six states (M6) at $h=1$ and $J_{k,[k+2]}=1.5$. (e) Dynamic phase diagram. The red circles, gray crosses, and magenta diamonds represent HC6, W3, and M6, respectively. The blue up-pointing triangles and green down-pointing triangles represent W6 and the intermediate wave mode (WI), respectively.
@@
Figure 2: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}=-1$ and $L=128$. (a) W6 at $h=1.5$. (b) WI at $h=1.2$. (c) HC6 at $h=0.8$. (c, inset) Enlarged density development at $t\sim 370~{}000$ showing skipping of $s=3$ and $4$ phases.
@@
Figure 3: Dynamics of the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$. (a–b) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states as functions of $h$ at (a) $J_{k,[k+2]}=-1$ and (b) $J_{k,[k+2]}=0$. The solid and dashed lines represent the data at $L=128$ and $256$, respectively. The bidirectional arrows at the top represent the ranges of dynamic modes at $J_{k,[k+2]}=-1$. (c) System-size dependence of phase boundaries. The solid and dashed lines represent the data at $J_{k,[k+2]}=-1$ and 00, respectively.
@@
Figure 4: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$, $h=1$, and $L=128$. (a) W3 at $J_{k,[k+2]}=0.5$. (b) Transition point at $J_{k,[k+2]}=1.214$. (c) M6 at $J_{k,[k+2]}=1.5$.
@@
This study focuses on the dynamics occurring under the cyclic symmetry, such that the flipping energies are constant as $h_{s,s^{\prime}}=h$ and the contact energy $J_{s,s^{\prime}}$ depends on the state distance $|s-s^{\prime}|$ only. A site is randomly selected and the flip to the neighboring states ($s=k\to[k+1]$ or $[k-1]$) is performed using the Metropolis Monte Carlo (MC) algorithm. This flip is performed $N$ times per MC step (time unit). We have checked that the choice of the Metropolis and Glauber rates for the update process introduces no qualitative change in the dynamics at $q=3$ [34].
@@
Through this study, $J_{k,k}=2$ and $J_{k,[k+1]}=0$, with $h$ and the other contact energies being varied: $J_{k,[k+2]}$ for $q=4$ and $5$, $J_{k,[k+2]}$ and $J_{k,[k+3]}$ for $q=6$, $J_{k,[k+2]}$, $J_{k,[k+3]}$, and $J_{k,[k+4]}$ for $q=8$. For the $q$-state standard Potts models in 2D under thermal equilibrium, the disorder-order transition is continuous and first-order for $q\leq 4$ and $q>4$, respectively, and the transition point is given by $J_{0,\mathrm{c}}=\ln(1+\sqrt{q})$ [16, 18]. Hence, the contact energy in this study, $J_{k,k}=2$, is sufficiently high for ordered phases in the standard Potts models to be obtained. The statistical errors were calculated from three or more independent runs (ten runs were used to calculate the scaling exponents).
@@ III.1.1 Modes for $J_{k,[k+3]}=J_{k,[k+2]}$
First, we describe the dynamics of the six-state Potts model ($q=6$) with $J_{k,[k+3]}=J_{k,[k+2]}$ (Figs. 16). Five modes are obtained as shown in Fig. 1.
[…]
@@
Figure 5: Dependence on $J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$, $h=1$, and $L=128$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$, $k$$[k+2]$, and $k$$[k+3]$ pairs.
@@
Figure 6: Transition between W3 and M6 modes at $J_{k,[k+3]}=J_{k,[k+2]}$ for (a–c) $h=0$ and (d–f) $h=1$. Binder cumulant $U_{3}$, susceptibility $\chi_{3}$, and order parameter $R_{3}$ for three-fold rotational symmetry are shown in (a,d), (b,e), and (c,f), respectively. The transition points are $J_{\mathrm{c}}=1.725$ and $1.214$ for $h=0$ and $1$, respectively.
@@
Figure 7: Six-state Potts model with $J_{k,[k+3]}=0$ and $L=128$. (a) Snapshot of homogeneous cycling of three states (HC3) at $h=1$ and $J_{k,[k+2]}=0.95$. (b) Snapshot of mixing of three states (M3) at $h=1$ and $J_{k,[k+2]}=1$. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC6, HC3, M3, W6, and WI, respectively.
@@
Figure 8: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=0$ and $h=1$. (a) HC3 at $J_{k,[k+2]}=0.976$. (b) M3 at $J_{k,[k+2]}=1$.
[…]
@@
For $J_{k,[k+2]}>0$, two additional modes emerge: wave modes of three states (W3) and a mixed phase of six states (M6), as shown in Figs. 1(c–e) and 4. As $J_{k,[k+2]}$ increases, the length of the boundary between the $s=k$ and $[k+2]$ domains increases owing to a reduction in the interfacial tension (see Fig. 5). When $J_{k,[k+2]}\simeq J_{k,k}=2$, the energy losses due to the contact between the different states (excluding the neighboring states $s=k$ and $[k+1]$) become negligibly small; hence, in the M6 mode, the six states are mixed to gain entropy. Note that this mode corresponds to the disordered phase at thermal equilibrium. For the intermediate $J_{k,[k+2]}$, three odd- or even-numbered states ($s=0,2,4$ or $s=1,3,5$) form domains exhibiting spiral waves, as observed in the three-state Potts model [34, 35] (see Fig. 1(c) and Movie S2). The other three states are present to small degrees at the domain boundaries. The boundary of the $s=k$ and $[k+2]$ domains moves via two-step flips ($s=k\to[k+1]\to[k+2]$), in which the clusters of $s=[k+1]$ states are swallowed up by the $[k+2]$ domain before growing. Waves of odd- and even-numbered states stochastically switch (see Fig. 4(a) and (b)). Note that this mode can be elucidated via symmetry factorization [55]. Static factorized phases for $q=2^{n}$ have been reported for the standard Ashkin-Teller model ($n=2$[21, 23, 24] and extended $n$-color models [22]. Here, the factorized (three-fold) symmetry remains within a $6=2\times 3$ symmetric system. For the wave propagation modes, the three-fold symmetry helps maintain their spiral centers [36].
@@
Figure 9: Dependence on $J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+3]}=0$ and $L=128$ for (a–b) $h=1$ and (c) $h=0.8$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$, $k$$[k+2]$, and $k$$[k+3]$ pairs. (c) Lifetimes of one-phase state ($\tau_{1}$) and HC3 mode ($\tau_{\mathrm{HC3}}$). The bidirectional arrows at the top represent the ranges of dynamic modes.
@@
The transition between the W3 and M6 modes is continuous. In the middle of the W3 range (Fig. 4(a)), the mean densities of the three dominant states are approximately $1/3$, and those of the other three states are approximately 00. With increasing $J_{k,[k+2]}$, the mean densities of the dominant and minor states decrease and increase, respectively. In the M6 mode, these mean densities converge to $1/6$ (see Fig. 4(b) and (c)). The transition between three and six states also occurs at thermal equilibrium ($h=0$), and the transition point $J_{k,[k+2]}=J_{\mathrm{c}}$ is weakly dependent on $h$ (see Fig. 1(e)). The transition is characterized by the order parameter $R_{3}$ for the three-fold symmetry. $R_{n}$ is defined as
@@
where $n$ is a positive integer and delimited-⟨$\langle$..$\rangle$ … ⟩ is the long-term time average. When all the sites are occupied by one state, $R_{n}=1$ for any $n$. When the sites are equally occupied by the three odd-numbered states, $R_{1}=0$ and $R_{3}=1$. When the sites are equally occupied by all six states, $R_{1}=R_{3}=0$. The susceptibility $\chi_{n}$ and Binder cumulant $U_{n}$ are defined, respectively, as [20, 39]

### v2

@@ Abstract
We studied the long-term nonequilibrium dynamics of $q$-state Potts models with $q=4$, 55, 66, and $8$ using Monte Carlo simulations on a two-dimensional square lattice. When the contact energies between the nearest neighbors for the standard Potts models are used, cyclic changes in the $q$ homogeneous phases and $q$-state coexisting wave mode appear at low and high flipping energies, respectively, for all values of $q$. However, for a factorizable $q$ value, dynamic modes with skipping states emerge, depending on the contact energies. For $q=6$, a spiral wave mode with three domain types (one state dominant or two states mixed) and cyclic changes in three homogeneous phases are found. Although three states can coexist spatially under thermal equilibrium, the scaling exponents of the transitions to the wave modes are modified from the equilibrium values.
@@ I Introduction
Various types of pattern formation have been observed in nonequilibrium systems [1, 2, 3, 4, 5, 6, 7]. Classically, they have been called self-organizing systems [1, 2], dissipative structures [1], and synergetics [2, 3], and recently, nonreciprocal systems [8, 9, 10, 11], featuring different aspects. In biological systems, spatiotemporal patterns have been observed from microscopic to macroscopic scales (cells and tissues to animal populations) [3, 7, 12, 13, 14, 15]. Under far-from-equilibrium conditions, the thermal fluctuations occurring in macroscopic patterns are typically negligible and the dynamics can be explained using deterministic continuum equations. In contrast, the thermal fluctuations can be significant in microscopic patterns, such as intracellular waves. However, the effects of the fluctuations and nucleation are not understood so far.
@@
Lattice models, including Ising and Potts models, have been widely used to study the phase transitions occurring at thermal equilibrium [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. The nonequilibrium dynamics have been investigated by imposing periodic external fields [31, 32] and spatial temperature inhomogeneity [33]. Recently, the cyclic flipping energies of three or four states have been introduced to these lattice models, and their pattern formations have been studied [34, 35, 36, 37, 38, 39]. We have used the standard Potts models [16, 17] for three [34, 35] and four [36] states in a two-dimensional (2D) square lattice under both cyclic symmetry [34, 36] and asymmetry [35, 36]; Hence, the transitions between dynamic modes, such as the cyclic changes of homogeneous phases and spiral waves, were revealed. These dynamics were also obtained in an off-lattice model for undulating fluid membranes [40]. These models exhibit stable spatiotemporal patterns in the long-term limit, unlike lattice Lotka–Volterra models, which exhibit an absorbing transition to a uniform state [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Manacorda and Fodor reported wave and global oscillation modes using a lattice model, which allows the occupancy of multiple particles at each site [37]. Further, Avni et al. have added a nonreciprocal interaction to the Ashkin-Teller model with a pairwise spin interaction (it is also called four-state vector-Potts and clock models [16]) and revealed changes in the scaling exponents of the second-order transition between disordered and ordered phases [38, 39]. However, this ordered phase maintains only for a short period ($t\propto$ system size $N$) starting from a uniform initial state.
@@
With the aim of clarifying the effects of thermal fluctuations on pattern formation and nonequilibrium phase transitions, this study investigates several transitions between the dynamic modes in the long-term limit ($t\to\infty$), which are independent of the initial states. We use general $q$-state Potts models with $q=4$, 55, 66, and $8$ under cyclically symmetric conditions. Various dynamic patterns are formed, particularly for a factorizable $q$ number. Although phases comprising factorized numbers of states have been reported at thermal equilibrium [21, 22, 23, 24, 55], to the best of our knowledge, spatiotemporal patterns with factorized symmetry have not been reported. We reveal several types of patterns with factorized symmetries and transitions between them.
@@
The model and methods are described in Sec. II. Simulation results are presented and discussed in Sec. III. The results for $q=6$, 55, 44, and $8$ are described in Secs. III.1,  III.2III.3, and III.4, respectively. The relationship with other models is discussed in Sec. IV. Finally, a summary is presented in Sec. V.
@@ II Active Potts Models
In standard Potts models, $J_{s_{i},s_{j}}=J_{0}\delta_{s_{i},s_{j}}$ is used. In equilibrium systems, each state can additionally have a self-energy $\varepsilon_{s}$, and the ratio of the forward and backward flip rates is $\exp(-\Delta H_{s_{i}s^{\prime}_{i}})$ for flipping of a single site from $s$ to $s^{\prime}$. The thermal energy $k_{\mathrm{B}}T$ is normalized to unity, $\Delta H_{s_{i}s^{\prime}_{i}}=\Delta H_{\mathrm{int}}-h_{s,s^{\prime}}$, and the flipping energy $h_{s,s^{\prime}}=\varepsilon_{s}-\varepsilon_{s}^{\prime}$. The cyclic sum of the flipping energy is $\sum_{k=0}^{q}h_{k,[k+1]}=0$ at equilibrium, where $[k]$ indicates $k\bmod q$. We extended this model to a nonequilibrium situation, in which $\sum_{k=0}^{q}h_{k,[k+1]}>0$, but $h_{ss^{\prime}}=-h_{s^{\prime}s}$ remains [34]. Hence, the detailed balance can be locally satisfied for flips between $s$ and $s^{\prime}$, but not globally for cycles ($s=0\to 1\ ...\to(q-1)\to 0$). For $q=3$, this corresponds to the rock–paper–scissors relationship. Note that this situation can be realized by reactions on a catalytic surface [56, 57, 58, 59, 60] and molecular transport through a membrane [61, 62, 63, 64, 40]. In the context of surface reactions, the state $s=0$ represents an unoccupied site, $s=1$ is a reactant-bound state, and $s=2$ to q2q-2 and s=q1s=q-1 are intermediate and final products, respectively. The bulk reaction energy for the final product is expressed as $\sum_{k=0}^{q}h_{k,[k+1]}$. In the case of molecular transport, the molecules have q1q-1 intermediate states in the membrane.
@@
Figure 1: Six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$ and $L=128$ (Pot6a). (a–d) Snapshots. The light yellow, light green, cyan, blue, magenta, and red sites (light to dark in grayscale) represent $s=0$, 11, 22, 33, 44, and $5$, respectively. (a) Waves of $s=3$ and $s=4$ domains in the intermediate wave mode (WI) at $h=1.2$ and $J_{k,[k+2]}=-1$. (b) Waves of six states (W6) at $h=1.5$ and $J_{k,[k+2]}=-1$. (c) Waves of three states (W3) at $h=1$ and $J_{k,[k+2]}=0.5$. (d) Mixing of six states (M6) at $h=1$ and $J_{k,[k+2]}=1.5$. (e) Dynamic phase diagram. The red circles, gray crosses, and magenta diamonds represent HC6, W3, and M6, respectively. The blue up-pointing triangles and green down-pointing triangles represent W6 and WI, respectively.
@@
Figure 2: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}=-1$ and $L=128$ (Pot6a). (a) W6 at $h=1.5$. (b) WI at $h=1.2$. (c) HC6 at $h=0.8$. (c, inset) Enlarged density development at $t\sim 370~000$ showing skipping of $s=3$ and 44 phases.
@@
Figure 3: Dynamic modes of the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}\leq 0$ (Pot6a). (a–b) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states as functions of $h$ at (a) $J_{k,[k+2]}=-1$ and (b) $J_{k,[k+2]}=0$. The solid and dashed lines represent the data at $L=128$ and $256$, respectively. The bidirectional arrows at the top represent the ranges of dynamic modes at $J_{k,[k+2]}=-1$. (c) System-size dependence of phase boundaries. The solid and dashed lines represent the data at $J_{k,[k+2]}=-1$ and 0, respectively. The statistical errors are shown as black bars. For most data points, they are smaller than the line thickness.
@@
Figure 4: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$, $h=1$, and $L=128$ (Pot6a). (a) W3 at $J_{k,[k+2]}=0.5$. (b) Transition point at $J_{k,[k+2]}=1.214$. (c) M6 at $J_{k,[k+2]}=1.5$.
@@
This study focuses on the dynamics occurring under the cyclic symmetry, such that the flipping energies are constant as $h_{s,s^{\prime}}=h$ and the contact energy $J_{s,s^{\prime}}$ depends on the state distance $|s-s^{\prime}|$ only. A site is randomly selected and the flip to the neighboring states ($s=k\to[k+1]$ or [k1][k-1]) is performed using the Metropolis Monte Carlo (MC) algorithm. This flip is performed $N$ times per MC step (time unit). We have checked that the choice of the Metropolis and Glauber rates for the update process introduces no qualitative change in the dynamics at $q=3$ [34].
@@
Through this study, $J_{k,k}=2$ and $J_{k,[k+1]}=0$, with $h$ and the other contact energies being varied: $J_{k,[k+2]}$ for $q=4$ and $5$, $J_{k,[k+2]}$ and $J_{k,[k+3]}$ for $q=6$, and $J_{k,[k+2]}$, $J_{k,[k+3]}$, and $J_{k,[k+4]}$ for $q=8$ ($J_{k,k^{\prime}}=J_{k^{\prime},k}$). For the $q$-state standard Potts models in 2D under thermal equilibrium, the disorder-order transition is continuous and first-order for $q\leq 4$ and $q>4$, respectively, and the transition point is given by $J_{0,\mathrm{c}}=\ln(1+\sqrt{q})$ [16, 18]. Hence, the contact energy in this study, $J_{k,k}=2$, is sufficiently high for ordered phases in the standard Potts models to be obtained. The statistical errors were calculated from three or more independent runs (ten runs were used to calculate the scaling exponents).
@@ III.1.1 Modes for $J_{k,[k+3]}=J_{k,[k+2]}$
For six-state Potts model ($q=6$), we have a 3D parameter space ($h,J_{k,[k+2]},J_{k,[k+3]}$), while keeping $J_{k,k}=2$ and $J_{k,[k+1]}=0$. We simulated the dynamics using parameters in three 2D slices ($J_{k,[k+3]}=J_{k,[k+2]}$, $J_{k,[k+3]}=0$, and $J_{k,[k+2]}=0$) to widely survey the parameter space. First, we describe the dynamics for $J_{k,[k+3]}=J_{k,[k+2]}$ (Figs. 16). We refer to this condition as Pot6a. Five modes are obtained as shown in Fig. 1.
For $J_{k,[k+2]}\leq 0$, negative contact energies are obtained only between neighbor pairs of the same states (i.e., repulsion between different states). Hence, a single state of $s\in[0,5]$ dominantly spans the entire lattice at $h=0$, and these six homogeneous phases are equivalent. At low but nonzero $h$, these six homogeneous phases change cyclically as $s=0\to 1\to 2\to 3\to 4\to 5\to 0$ (see Fig. 2(c)). This is a homogeneous cycling mode, as previously observed in the three- and four-state active standard Potts models [34, 36]. In this paper, this mode is abbreviated as HC6 (that is, HCnn denotes the cycling of $n$ states). Note that cyclic changes occur through nucleation and growth, and the nucleation period decreases exponentially with increasing $h$ [34]. When a nucleus of the next state ($s=[k+1]$) appears in the growth of an $s=k$ domain, the $s=k$ dominant phase is skipped [36]. Skipping of the $s=3$ and $4$ phases is shown in Fig. 2(c, inset). As $h$ increases, the subsequent nuclei form more frequently during domain growth (Fig. 1(a)); subsequently, spatial coexistence of multiple states becomes dominant. At high $h$, all six states spatially coexist, and the domain boundary between neighboring states $s=k$ and $[k+1]$ moves ballistically in the direction from the $s=[k+1]$ to $s=k$ domains (see Fig. 2(a) and Movie S1). Herein, this mode is abbreviated as W6 (where Wnn denotes the waves of $n$ states). With increasing repulsion $|J_{k,[k+2]}|$, the boundary length between the $s=k$ and $[k+2]$ or $[k+3]$ domains decreases, such that the waves more frequently form a spiral shape (see Movie S1). At intermediate values of $h$, intermediate numbers of states (2255) spatially coexist for most of the period (see the bottom region in Fig. 2(b). A few states remain at $N_{s}/N\simeq 0$ for most of the period). In this study, this intermediate wave mode is abbreviated as WI. This behavior differs from those in three- and four-state active standard Potts models [34, 36], in which HCqq and Wqq modes temporally coexist in the intermediate conditions. For $q=6$, HC6 and W6 modes temporally appear in the WI mode but only for short periods, such that the lattice is primarily occupied by intermediate numbers of states.
@@
Figure 5: Dependence on $J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+3]}=J_{k,[k+2]}$, $h=1$, and $L=128$ (Pot6a). (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for kk[k+1][k+1], kk[k+2][k+2], and kk[k+3][k+3] pairs.
@@
Figure 6: Transition between W3 (M3) and M6 modes at $J_{k,[k+3]}=J_{k,[k+2]}$ for (a–c) $h=0$ and (d–f) $h=1$ (Pot6a). Binder cumulant $U_{3}$, susceptibility $\chi_{3}$, and order parameter $R_{3}$ for three-fold rotational symmetry are shown in (a,d), (b,e), and (c,f), respectively. The transition points are $J_{\mathrm{c}}=1.725$ and $1.214$ for $h=0$ and $1$, respectively.
@@
Figure 7: Six-state Potts model with $J_{k,[k+3]}=0$ and $L=128$ (Pot6b). (a) Snapshot of homogeneous cycling of three states (HC3) at $h=1$ and $J_{k,[k+2]}=0.95$. (b) Snapshot of mixing of three states (M3) at $h=1$ and $J_{k,[k+2]}=1$. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC6, HC3, M3, W6, and WI, respectively.
@@
Figure 8: Time development of number densities of states in the six-state Potts model at $J_{k,[k+3]}=0$ and $h=1$ (Pot6b). (a) HC3 at $J_{k,[k+2]}=0.976$. (b) M3 at $J_{k,[k+2]}=1$.
To distinguish between these three modes, we calculate the time fractions $p_{\mathrm{phase}}$ of single phases and multi-phase coexistence. It is considered that the lattice is covered by one phase at $N_{s}/N>0.95$ for $s\in[0,q-1]$, and that $n$ phases spatially coexist when $n$ states satisfy $N_{s}/N>0.05$. The fractions of the one-phase dominance and six-phase coexistence decreases and increases with increasing $h$, respectively (see Fig. 3(a) and (b)). For the HC and W6 modes, the largest fractions are obtained for the one-phase dominance and six-phase coexistence, respectively. For the WI mode, the largest fraction is obtained for one of the $n$-phase coexistences, where n=2n=255 (see the bidirectional arrows at the top of Fig. 3(a)). As the system size increases, the transition points of $h$ between HC6 and WI and between WI and W6 decrease (see Fig. 3), similar to those between HCqq and Wqq in the three- and four-state active standard Potts models [34, 36].
@@
For $J_{k,[k+2]}>0$, two additional modes emerge: wave modes of three states (W3) and a mixed phase of six states (M6), as shown in Figs. 1(c–e) and 4. As $J_{k,[k+2]}$ increases, the length of the boundary between the $s=k$ and $[k+2]$ domains increases owing to a reduction in the interfacial tension (see Fig. 5). When $J_{k,[k+2]}\simeq J_{k,k}=2$, the energy losses due to the contact between the different states (excluding the neighboring states $s=k$ and $[k+1]$) become negligibly small; hence, in the M6 mode, the six states are mixed to gain entropy. Note that this mode corresponds to the disordered phase at thermal equilibrium. For the intermediate $J_{k,[k+2]}$, three odd- or even-numbered states ($s=1,3,5$ or $s=0,2,4$) form domains exhibiting spiral waves, as observed in the three-state Potts model [34, 35] (see Fig. 1(c) and Movie S2). The other three states are present to small degrees at the domain boundaries. The boundary of the $s=k$ and $[k+2]$ domains moves via two-step flips ($s=k\to[k+1]\to[k+2]$), in which the clusters of $s=[k+1]$ states are swallowed up by the $[k+2]$ domain before growing. Waves of odd- and even-numbered states stochastically switch (see Fig. 4(a) and (b)). Note that this mode can be elucidated via symmetry factorization [55]. Static factorized phases for $q=2^{n}$ have been reported for the standard Ashkin-Teller model (n=2n=2[21, 23, 24] and extended $n$-color models [22]. Here, the factorized (three-fold) symmetry remains within a $6=2\times 3$ symmetric system. For the wave propagation modes, the three-fold symmetry helps maintain their spiral centers [36].
@@
Figure 9: Dependence on $J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+3]}=0$ and $L=128$ for (a–b) $h=1$ and (c) $h=0.8$ Pot6b). (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. (b) Contact probabilities $p_{\mathrm{contact}}$ for kk[k+1][k+1], kk[k+2][k+2], and kk[k+3][k+3] pairs. (c) Lifetimes of one-phase state ($\tau_{1}$) and HC3 mode ($\tau_{\mathrm{HC3}}$). The bidirectional arrows at the top represent the ranges of dynamic modes.
@@
The transition between the W3 and M6 modes is continuous. In the middle of the W3 range (Fig. 4(a)), the mean densities of the three dominant states are approximately $1/3$, and those of the other three states are approximately 0. With increasing $J_{k,[k+2]}$, the mean densities of the dominant and minor states decrease and increase, respectively. In the M6 mode, these mean densities converge to $1/6$ (see Fig. 4(b) and (c)). The transition between three and six states (M3 and M6) also occurs at thermal equilibrium ($h=0$), and the transition point $J_{k,[k+2]}=J_{\mathrm{c}}$ is weakly dependent on $h$ (see Fig. 1(e)). The transition is characterized by the order parameter $R_{3}$ for the three-fold symmetry. $R_{n}$ is defined as
@@
where $n$ is a positive integer and $\langle$$\rangle$ is the long-term time average. When all the sites are occupied by one state, $R_{n}=1$ for any $n$. When the sites are equally occupied by the three odd-numbered states, $R_{1}=0$ and $R_{3}=1$. When the sites are equally occupied by all six states, $R_{1}=R_{3}=0$. The susceptibility $\chi_{n}$ and Binder cumulant $U_{n}$ are defined, respectively, as [20, 39]